OpenStudy (anonymous):

Explain how you would find the exact value of 3 over the square root of 8?

OpenStudy (anonymous):



OpenStudy (anonymous):

You will need to first simplify the square root of 8. Once you get this you will have something like the following \[ \frac{3}{x\sqrt{y}} \] Form here you need to realise that you can multiply this by a form of one in the form of sqrt(y)/sqrt(y) so you get \[ \frac{\sqrt{y}}{\sqrt{y}}\frac{3}{x\sqrt{y}} = \frac{3\sqrt{y}}{xy}\] that will give you the simplified exact answer.

OpenStudy (anonymous):

oh okay. Thanks!

OpenStudy (anonymous):

So whats the simplified answer?

OpenStudy (anonymous):

WE cannt give the answer to you but we can help you. Can you show work or do you know how to simplify sqrt(8) ?

OpenStudy (anonymous):

I dont know how

OpenStudy (anonymous):

So you need to find out what multiplies by what to get 8. So you have 1 and 8. 2 and 4, and that is it. From there you can use the following property to split it up. \[ \sqrt{x^2 * y} = x\sqrt{y} \] So if you do that, what do you get?

OpenStudy (anonymous):

but there are no numbers

OpenStudy (anonymous):

the x and y represent numbers. You have to plug them into each other. But I guess I am getting a bit ahead of you as you may not have taken algebra yet (btw a programing language greatly helps with getting good with those) So you have \[ \sqrt(8) = \sqrt{4 * 2} \] this can be written as \[ \sqrt{2^2 * 2} \] which is very simalar to the equation I put up before right? That is equal to \[ 2\sqrt{2} \] So you need to replace the sqrt(8) in the equation with that and then multiply by a form of one as I said in the first post (with y = 2 or where you see a y you write a 2)

OpenStudy (anonymous):

ohh okay.